everettdrums:

Which of the following expressions is equivalent to (y − 4)4? y4 − 48y3 + 8y2 − 4 y3 − 12y2 + 48y − 64 y4 − 16y3 + 96y2 − 256y + 256 y4 + 16y3 + 96y2 + 256y + 256

FortyTheMC:

So (y-4)4 is the same as (y-4) but 4 times. So write it out as (y-4)(y-4)(y-4)(y-4) and start distributing and see how far you can get

jhonyy9:

\[\left( x - 4 \right)^{4} = \left( x - 4 \right)^{2} × \left( x - 4 \right)^{2} = ?\] try continue from this